OpenStudy (anonymous):

I need help, when (3x^3 + px^2 + qx +8) is divided by (x^2 - 3x +2) , the remainder is (5x+6). how do we find the values of p and q? thank you :)

myininaya (myininaya):

i will scan this and post it here too long to do by latex

myininaya (myininaya):

i just did long division and now we have a set of linear equations to solve

OpenStudy (campbell_st):

the divisor x^2 -3x +2 is factorised to (x-2)(x-1) so you can find P(2) and P(1)

myininaya (myininaya):

so we have from what I did the following equations: 3p+q=-16 ; -2p-10=6 -2p=16 p=-8 now just plug this into our other equation to find q 3(-8)+q=-16 -24+q=-16

OpenStudy (anonymous):

thanks for taking the trouble to help me solve this question :) i rly appreciate it!

OpenStudy (anonymous):

thank you guys! :)